OpenStudy (princesssleelee):

Three functions are given below: f(x), g(x), and h(x). Explain how to find the axis of symmetry for each function, and rank the functions based on their axis of symmetry (from smallest to largest).

OpenStudy (princesssleelee):

f(x) = –4(x − 8)2 + 3 g(x) = 3x2 + 12x + 15

OpenStudy (princesssleelee):

axis of symmetry=vertex :)

OpenStudy (mathmale):

Please take care to express "exponentiation" as (x-8)^2 [not as (x-8)2 ] and 3x^2 [not as 3x2]. Both f(x) and g(x), if simplified, will appear in the form of a quadratic function. Please do this first. Your results should have the form y=ax^2 + bx + c. The equation of the axis of symmetry is x=-b/(2a).

OpenStudy (princesssleelee):

@mathmale i dont understand

OpenStudy (mathmale):

What part was not clear for you? Focus on rewriting your two functions first. Your results should be in the form y=ax^2 + bx + c, or, in other words, in descending order by powers of x. highest power: x^2 ("x squared") next power: x Next: the constant, c